OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Show each step of your process.

OpenStudy (anonymous):

I think I found the answer but don't know the process.

OpenStudy (anonymous):

@kevindragonlord How can you get the answer w/o some form of process?

OpenStudy (anonymous):

internet

OpenStudy (anonymous):

I'm trying to learn the process right now

OpenStudy (anonymous):

Thank you for not providing the direct answer :)

OpenStudy (anonymous):

Ok. :)

OpenStudy (anonymous):

@Aureyliant can you help?

OpenStudy (anonymous):

\[\sqrt{x}*\sqrt{x}=x ^{\frac{ 1 }{ 2 }}*x ^{\frac{ 1 }{ 2 }}=x ^{\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }}=x ^{\frac{ 1+1 }{ 2 } }=?\]

OpenStudy (anonymous):

x^2/2

OpenStudy (anonymous):

@surjithayer hello?

OpenStudy (anonymous):

@animalfan1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

\[x ^{\frac{ 2 }{ 2 }}=x^1=x\]